Guys please help me with this problem, thanks in advanced! 200 mg of KOH are added with H2O until a final volume of 720 ml is reached. Determine the pH of the solution obtained and tell if the solution is acidic or basic.
1
Expert's answer
2018-04-16T11:02:08-0400
pH = -lg [H+]
СM = n/V
n = m/M
pH + pOH = 14
M (KOH) = 56.1 g/mol
n (KOH) = 0.2/56.1 = 0.004 mol
СM = 0.004 / 0.72 = 0.005 mol/l
pOH = - lg [0.005] = 2.3
pH = 14 – 2.3 = 11.7
The solution will be basic.